# On the importance of parallel magnetic-field fluctuations for electromagnetic instabilities in STEP

### v1

@@ 6 The importance of parallel magnetic perturbations to the hKBM
In [9] it was shown that the total heat and particle fluxes from nonlinear simulations vary with $\beta$ when $\beta^{\prime}$ is varied consistently (see Fig. 11 of [9]). However, the heat and particle fluxes remained very large even at much smaller values of $\beta$ than the STEP baseline. It was argued that reduced $\beta^{\prime}$ stabilisation of the hKBM plays a significant role in maintaining such large turbulent fluxes. In the STEP baseline case, $\beta^{\prime}$ stabilisation is very important since it pushes the hKBM towards marginality. The sensitive dependence of the hKBM on $\beta^{\prime}$ is made most obvious when $\beta^{\prime}$ is varied holding all other local equilibrium parameters fixed (albeit inconsistent).
Fig 8 shows the linear growth rate (a, c) and frequency (b, d) as functions of the binormal wavenumber for simulations with reduced $\beta^{\prime}$ with, $f_{B}=1$ (blue), and without, $f_{B}=0$ (orange, green, red), $\delta\!B_{\parallel}$ fluctuations. In all cases the hKBM is much more unstable than the case with the nominal value of $\beta^{\prime}$ (see Fig. 2) due to the much decreased role of $\beta^{\prime}$ stabilisation. In each case, the hKBM is likely to drive much more transport (note also that the mode does not become stable at $n=1$ and this is certainly a no-go zone for any actual device).
There are three key observations from Fig 8 that we wish to stress.
  • In the limit $\beta^{\prime}\rightarrow 0,$ the MHD-1 and MHD-2 are identical, which is not surprising as $\mathbf{b}_{0}\cdot\mathbf{\nabla B_{0}}\rightarrow\mathbf{\nabla}B_{0}/B_{0}$ as $\mathbf{\nabla}p\rightarrow 0$ (see C). This is clearly seen in Fig 8 where the difference between the MHD-1 and MHD-2 prediction reduces as $|\beta^{\prime}|$ decreases.
  • We note that the MHD approximation does a much better job of capturing the linear growth rates of the hKBM at smaller values of $|\beta^{\prime}|$. The limit $|\beta^{\prime}|\rightarrow 0$ is in effect a $\omega_{\nabla B}-\omega_{\kappa}\rightarrow 0$ approximation but this does not say anything about $\delta\!B_{\parallel}.$
  • The hKBM can be unstable up to $n=1$ in the absence of $\delta\!B_{\parallel}$ with exact drifts.
@@
Figure 9: Time traces of the total heat flux (a) and total particle flux (b) from GENE simulations. Simulations are shown both with, $f_{B}=1$ (blue), and without, $f_{B}=0$ (magenta), $\delta B_{\parallel}$. In the case without $\delta\!B_{\parallel},$ $\beta^{\prime}=0.4\beta^{\prime}_{\mathrm{STEP}}.$ No approximations are made concerning the magnetic-drift velocity.
This final result suggests it should be possible to study hKBM-driven turbulence in more strongly driven regimes using global GK codes that do not include $B_{\parallel}$.
Fig 9 shows time traces of the total heat flux and particle flux from two nonlinear GENE simulations: (i) with the nominal value of $\beta^{\prime}$ that includes $\delta\!B_{\parallel}$ fluctuations (blue); and (ii) excluding $\delta\!B_{\parallel}$ fluctuations and using exact drifts (magenta) at a lower value of $\beta^{\prime}=0.4\beta^{\prime}_{\mathrm{STEP}}$ (with all other parameters are identical). As expected from the linear spectrum (Fig 8), the nonlinear physics of this new case qualitatively mirrors the STEP baseline case even without $\delta\!B_{\parallel}$ fluctuations. The heat fluxes rise to very large levels and no robustly steady saturated state is reached in this simulation on the timescale simulated (note that this is the only simulation in this paper that does not reach a robustly steady saturated state). Snapshots of the turbulent fields Fig 10(g,h,i) reveal the presence of box-scale radially elongated streamers.
The fact that at lower $\beta^{\prime}$ we obtain electromagnetic runaway fluxes without $\delta\!B_{\parallel}$ is an important result; it reduces the number of ingredients needed to observe electromagnetic runaway fluxes in local gyrokinetic simulations. Put more simply, we observe qualitatively similar runaway fluxes in local simulations when there is electromagnetic instability at the longest-wavelengths in the system. We take great care to emphasise that the results in this paper suggests that there is an important distinction to be made between (A) runaway fluxes that eventually saturate at very large values such as those discussed in this paper, and (B) fluxes that will never saturate such as the runaways observed by [31] in cyclone-base-case[32] geometry without $\delta\!B_{\parallel}.$ We find that results for the STEP equilibrium considered in this paper definitively fall into category (A). It is left unclear which category the results at lower $\beta^{\prime}$ fall into. Identification of these two categories motivates the need for a rigorous theory of electromagnetic turbulence saturation that can distinguish between the two types of behaviour.
@@ 7.1 The MHD approximation is typically not appropriate for use in high-$\beta_{e}$ reactor-scale plasmas
In Section 4, it was shown that inclusion of $\delta\!B_{\parallel}$ and a rigorous treatment of the magnetic drift velocity is essential to capture the correct linear spectrum of the hKBM in the STEP flat-top. The MHD approximation [^13] The MHD approximation is most faithfully represented by variant MHD-1, discussed in the main text, which compensates for the neglect of parallel magnetic perturbations by setting $\omega_{\nabla B}=\omega_{\kappa}$. roughly captures the linear spectrum of the hKBM, but predicts lower growth rates in the STEP local equilibrium and is significantly more stable at long wavelengths.
@@
Even though the MHD approximation (MHD-1) reproduces qualitatively similar linear physics, Section 5 reveals that the corresponding nonlinear simulations do not accurately track nonlinear simulations for STEP with full physics that were reported in [9]. Typically, simulations which attempt to resolve the hKBM with full physics (including $\delta\!B_{\parallel}$) yield fluxes rising to very large values and it can be difficult to obtain a robustly steady saturation period in the absence of flow shear (i.e. this typically requires a large amount of computational resource). When the MHD-1 approximation is used, the turbulence saturates at a level around two orders of magnitude lower than that reached in a comparable time in the full physics simulations. The reason for this difference is that the longest wavelength (i.e., modes corresponding to $n<10$) are stable when the MHD approximation is used, but unstable with full physics. It was argued that these linearly unstable modes at very long-wavelength are responsible for the transition to very large fluxes (see Fig 7 and discussion).
In Section 4.2 it was found very difficult to improve the linear performance of the MHD-approximation whilst making only relatively minor changes to the local equilibrium.
@@
7.2 $\delta\!B_{\parallel}$ is not essential for studying hKBM driven turbulence. However, $\delta\!B_{\parallel}$ is essential for studying turbulence in the high $\beta^{\prime}$ conditions of the STEP flat-top.
@@
In Section 6, $\beta^{\prime}$ stabilisation [30] was identified as being responsible for the sensitivity of hKBM stability to $\delta\!B_{\parallel}$ fluctuations, and it was noted that this plasma sits close to marginality. It was also shown, however, that the hybrid mode becomes unstable in the absence of $\delta\!B_{\parallel}$ if $\beta^{\prime}$ is reduced to push the local equilibrium that are further from marginality. In such local equilibria (further from marginal stability at long wavelength), electromagnetic runaway fluxes emerge in nonlinear simulations that neglect $\delta\!B_{\parallel}$ (with or without the MHD approximation). This suggest that turbulence can be simulated in some high-$\beta$ ST plasmas using presently available tools and codes that neglect $B_{\parallel}$.
@@
In the present design of the STEP flat-top, however, it is essential to include $B_{\parallel}$ because the modes expected to dominate turbulence are both: (i) unstable only with $\delta\!B_{\parallel}$ and; (ii) unstable up very long wavelengths. Global gyrokinetic codes currently being developed to include $\delta\!B_{\parallel}$ rigorously will be important and timely for STEP.
@@ Appendix A The MHD approximation in the GK framework
×$\times\left[\delta\!\phi-v_{\parallel}\delta\!A_{\parallel}+\frac{m_{s}v_{\perp}^{2}}{2q_{s}B_{0}}\delta\!B_{\parallel}\right]F_{0s}.$
@@
$\delta P_{\perp}=\sum_{s}\int\mathrm{d}\mathbf{v}^{3}\,F_{0s}\frac{m_{s}v_{\perp}^{2}q_{s}}{2T_{0s}}\left[1-\frac{\omega_{\star s}^{T}}{\omega}+\frac{\omega_{ds}}{\omega}\right]\times\left[\delta\!\phi-\frac{m_{s}v_{\perp}^{2}}{2q_{s}}\frac{\mu_{0}\delta P_{\perp}}{B_{0}^{2}}\right].$ (15)
@@
$\delta\!P_{\perp}$ $=$ $\sum_{s}\int\mathrm{d}\mathbf{v}^{3}\,F_{0s}\left\{\frac{m_{s}v_{\perp}^{2}}{2T_{0s}}q_{s}\delta\!\phi\left[1-\frac{\omega_{\star s}^{T}}{\omega}+\frac{\omega_{ds}}{\omega}\right]\right.-\left.\frac{m_{s}^{2}v_{\perp}^{4}}{4T_{0s}}\frac{\mu_{0}\delta P_{\perp}}{B_{0}^{2}}\left[1-\frac{\omega_{\star s}^{T}}{\omega}+\frac{\omega_{ds}}{\omega}\right]\right\}$
$=$ $\underbrace{\delta\!\phi\sum_{s}\int\mathrm{d}\mathbf{v}^{3}\,\frac{q_{s}m_{s}v_{\perp}^{2}F_{0s}}{2T_{0s}}\left[1-\frac{\omega_{\star s}^{T}}{\omega}+\frac{\omega_{ds}}{\omega}\right]}_{[I]}-\underbrace{\frac{\mu_{0}\delta\!P_{\perp}}{B_{0}^{2}}\sum_{s}\int\mathrm{d}\mathbf{v}^{3}\,\frac{m_{s}^{2}v_{\perp}^{4}F_{0s}}{4T_{0s}}\left[1-\frac{\omega_{\star s}^{T}}{\omega}+\frac{\omega_{ds}}{\omega}\right]}_{[II]}$
@@
$[I]$ $=$ $\delta\!\phi\sum_{s}n_{s}q_{s}\left\{1-\frac{\omega_{\star s}}{\omega}\left(1+\eta_{s}\right)+\frac{T_{0s}}{q_{s}B_{0}\omega}\left(\omega_{\kappa}+2\omega_{\nabla B}\right)\right\}.$ (17)
The first term in (17) vanishes due to quasineutrality. On summing over species the second term reduces to:
$\sum_{s}n_{s}q_{s}\omega_{\star s}(1+\eta_{s})=\frac{(\mathbf{k}\times\mathbf{b}_{0})\nabla{P}}{B_{0}}=\frac{B_{0}}{\mu_{0}}\left(\omega_{\kappa}-\omega_{\nabla B}\right),$ (18)
@@
$[I]$ $=\frac{B_{0}\delta\!\phi}{\mu_{0}\omega}\left(\left(\omega_{\nabla B}-\omega_{\kappa}\right)+\frac{\beta}{2}\left(\omega_{\kappa}+2\omega_{\nabla B}\right)\right)$ (19)
@@
$[II]=-\delta P_{\perp}\sum_{s}\beta_{s}\left[1-\frac{\omega_{\star s}}{\omega}(1+2\eta_{s})+\frac{T_{0s}}{q_{s}B_{0}\omega}(\omega_{\kappa}+3\omega_{\nabla B})\right]$ (20)
@@
$\delta P_{\perp}=\frac{B_{0}\delta\!\phi}{\mu_{0}\omega}\left(\omega_{\nabla B}-\omega_{\kappa}\right),$ (21)
@@
Appendix B Cancellation of Magnetic Drift and $\mathbf{B_{\parallel}}$ Terms in Appropriate Limits
Using the relationship between $\nabla B$ and curvature drifts in 32, the drift term on the left-hand side of the GK equation in 1 can be written in a form convenient for the MHD-1 approximation:
$T_{ds}$ $=$ $-\frac{1}{\Omega_{s}}\left[\omega_{\kappa}\left(v_{\parallel}^{2}+\frac{v_{\perp}^{2}}{2}\right)-\frac{\mu_{0}(\mathbf{k}_{\perp}\cdot\mathbf{b})\nabla p}{B_{0}^{2}}\frac{v_{\perp}^{2}}{2}\right]g_{s}$ (23)

### v2

@@ 6 The importance of parallel magnetic perturbations to the hKBM
In [9] it was shown that the total heat and particle fluxes from nonlinear simulations vary with $\beta$ when $\beta^{\prime}$ is varied consistently (see Fig. 11 of [9]). However, the heat and particle fluxes remained very large even at much smaller values of $\beta$ than the STEP baseline. It was argued that reduced $\beta^{\prime}$ stabilisation of the hKBM is largely responsible for such large turbulent fluxes at lower $\beta$. In the STEP baseline case, $\beta^{\prime}$ stabilisation is very important since it pushes the hKBM towards marginality. The sensitive dependence of the hKBM on $\beta^{\prime}$ is made most obvious when $\beta^{\prime}$ is varied holding all other local equilibrium parameters fixed (albeit inconsistent). It is interesting to note, as remarked in Section 4.2.2, that we should expect the MHD approximation to improve as $|\beta^{\prime}|$ is reduced.
Fig 6 shows the linear growth rate (a, c) and frequency (b, d) as functions of the binormal wavenumber for simulations with reduced $\beta^{\prime}$ with, $f_{B}=1$ (blue), and without, $f_{B}=0$ (orange, green, red), $\delta\!B_{\parallel}$ fluctuations. In all cases the hKBM is much more unstable than at the nominal value of $\beta^{\prime}$ (see Fig. 2) due to the much decreased role of $\beta^{\prime}$ stabilisation. In each case, the hKBM is likely to drive much more transport.
There are several key observations from Fig 6 that we wish to stress.
  • In the limit $\beta^{\prime}\rightarrow 0,$ the MHD-1 and MHD-2 are identical, which is not surprising as $\mathbf{b}\cdot\mathbf{\nabla}\mathbf{b}\rightarrow(\mathbf{\nabla_{\perp}}B)/B$ as $\mathbf{\nabla}p\rightarrow 0$ (see C). This is clearly seen in Fig 6 where the difference between the MHD-1 and MHD-2 prediction reduces as $|\beta^{\prime}|$ decreases. The limit $|\beta^{\prime}|\rightarrow 0$ is in effect a $\omega_{\nabla B}-\omega_{\kappa}\rightarrow 0$ approximation. In this physical limit $\delta\!B_{\parallel}=0$, and the MHD approximation should be trivially satisfied.
  • We note that the MHD approximation does a much better job of capturing the linear growth rates of the hKBM at smaller values of $|\beta^{\prime}|$, in agreement with the discussion in Section 4.2.2.
  • The hKBM can be unstable up to $n=1$ in the absence of $\delta\!B_{\parallel}$ with exact drifts, and this is certainly a no-go zone for any actual device.
In more strongly driven regimes at lower $\beta^{\prime}$ in STEP, it is clearly possible to study hKBM-driven turbulence using global GK codes that do not include $B_{\parallel}$.
@@
Figure 7: Time traces of the total heat flux (a) and total particle flux (b) from GENE simulations. Simulations are shown both with $\delta B_{\parallel}$ at the nominal $\beta^{\prime}$ ($f_{B}=1$, blue), and without at $\beta^{\prime}=0.4\beta^{\prime}_{\mathrm{STEP}}$ ($f_{B}=0$, magenta). No approximations are made concerning the magnetic-drift velocity. The lower $\beta^{\prime}$ simulation without $\delta\!B_{\parallel}$ (magenta) exhibits a qualitatively similar electromagnetic runaway (transition to very large heat fluxes) to that seen with full physics since the hKBM has been driven unstable at long wavelength, but it is the only simulation in this paper that does not meet the statistical definition of saturation introduced in [9].
Fig 7 shows time traces of the total heat flux and particle flux from two nonlinear GENE simulations: (i) with the nominal value of $\beta^{\prime}$ that includes $\delta\!B_{\parallel}$ fluctuations (blue); and (ii) excluding $\delta\!B_{\parallel}$ fluctuations and using exact drifts (magenta) at a lower value of $\beta^{\prime}=0.4\beta^{\prime}_{\mathrm{STEP}}$ (with all other parameters are identical). As expected from the linear spectrum (Fig 6), the nonlinear physics of this new case qualitatively mirrors the STEP baseline case even without $\delta\!B_{\parallel}$ fluctuations. The heat fluxes rise to very large levels and no robustly steady saturated state is reached over the period of this simulation: (note that this is the only simulation in this paper that does not reach a robustly steady saturated state). It is currently unknown whether this simulation would obtain a steady state if it were continued and such a question is very difficult to answer numerically due to the computational cost. Simulations such as those discussed here motivate the pressing need for a rigorous theory of electromagnetic turbulence saturation in the flux-tube limit. Snapshots of the turbulent fields Fig 8(g,h,i) reveal the presence of box-scale radially elongated streamers.
The fact that at lower $\beta^{\prime}$ we obtain electromagnetic runaway fluxes without $\delta\!B_{\parallel}$ is an important result; it reduces the number of ingredients needed to observe electromagnetic runaway fluxes in local gyrokinetic simulations. Put more simply, we observe qualitatively similar runaway fluxes in local simulations when there is electromagnetic instability at the longest-wavelengths in the system. We take great care to emphasise that the results in this paper suggests that there is an important distinction to be made between (A) runaway fluxes that eventually saturate at very large values such as those discussed in this paper, and (B) fluxes that never saturate such as the runaways reported by [33] for cyclone-base-case[34] geometry without $\delta\!B_{\parallel}.$ All nonlinear results for the STEP equilibrium at the nominal $\beta^{\prime}$ in this paper definitively fall into category (A). It remains unclear which category the lower $\beta^{\prime}$ simulation falls into. Identification of these two categories motivates the need for a rigorous theory of electromagnetic turbulence saturation that can distinguish between the two types of behaviour.
@@ 7.1 The MHD approximation is typically not appropriate for use in high-$\beta_{e}$ reactor-scale plasmas
In Section 4.2.2 it was argued that the MHD approximation [^11] The MHD approximation is faithfully represented by the implementation MHD-1, discussed in the main text, which compensates for the neglect of parallel magnetic perturbations by setting $\omega_{\nabla B}=\omega_{\kappa}$. should perform well in the limit $\beta^{\prime}\ll\sqrt{\beta}\ll 1$. This limit is not well satisfied in STEP-EC-HD (where $\beta_{e}=0.09$ and $|\beta^{\prime}|=0.48$ at $\rho/a=0.64$), and it may also be questionable in existing experiments at internal/external transport barriers with high pressure gradients, and in other reactor-scale plasmas.
@@
In Section 4, it was shown that inclusion of $\delta\!B_{\parallel}$ and a rigorous treatment of the magnetic drift velocity is essential to capture the correct linear spectrum of the hKBM in the STEP flat-top. The MHD approximation roughly captures the linear spectrum of the hKBM, but predicts lower growth rates and is significantly more stable at long wavelengths for this STEP local equilibrium.
@@
Even though the MHD approximation reproduces qualitatively similar linear physics, Section 5 reveals that the corresponding nonlinear simulations do not accurately track nonlinear simulations for STEP with full physics that were reported in [9]. Typically, simulations which attempt to resolve the hKBM with full physics (including $\delta\!B_{\parallel}$) yield fluxes rising to very large values and require substantial computational resource to reach saturation. When the MHD approximation is used, the turbulence saturates at a level around two orders of magnitude lower than that reached in a comparable time in the full physics simulations. The reason for this difference is that the longest wavelength modes (i.e., corresponding to $n<10$) are stable under the MHD approximation, but unstable with full physics. It was argued that these linearly unstable modes at very long-wavelength are responsible for the transition to very large fluxes (see Fig 5 and discussion).
@@
7.2 $\delta\!B_{\parallel}$ is not always essential for studying hKBM driven turbulence. However, $\delta\!B_{\parallel}$ is essential for studying turbulence in the high $\beta^{\prime}$ conditions of the STEP flat-top.
@@
In Section 6, strong $\beta^{\prime}$ stabilisation [32] was identified as being responsible for the sensitivity of hKBM stability to $\delta\!B_{\parallel}$ fluctuations, and it was noted that this plasma sits close to marginality. It was also shown, however, that the hybrid mode becomes unstable in the absence of $\delta\!B_{\parallel}$ if $\beta^{\prime}$ is reduced to push the local equilibrium further from marginality. In such local equilibria, which are more unstable at long wavelength, electromagnetic runaway fluxes emerge in nonlinear simulations that neglect $\delta\!B_{\parallel}$ (without, and likely also with, the MHD approximation). This suggest that turbulence can be simulated in some high-$\beta$ ST plasmas using presently available tools and codes that neglect $\delta\!B_{\parallel}$.
@@
In the present design of the STEP flat-top, however, it is essential to include $\delta\!B_{\parallel}$ because the modes expected to dominate turbulence are both: (i) unstable only with $\delta\!B_{\parallel}$ and; (ii) unstable up to very long wavelengths. Global gyrokinetic codes currently being developed to include $\delta\!B_{\parallel}$ rigorously will be important and timely for STEP.
@@ Appendix A The MHD approximation in the GK framework
×$\times\left[\delta\!\phi-v_{\parallel}\delta\!A_{\parallel}+\frac{m_{s}v_{\perp}^{2}}{2q_{s}B}\delta\!B_{\parallel}\right]F_{0s}.$
@@
$\delta P_{\perp}=\sum_{s}\int\mathrm{d}\mathbf{v}^{3}\,F_{0s}\frac{m_{s}v_{\perp}^{2}q_{s}}{2T_{0s}}\left[1-\frac{\omega_{\star s}^{T}}{\omega}+\frac{\omega_{ds}}{\omega}\right]\times\left[\delta\!\phi-\frac{m_{s}v_{\perp}^{2}}{2q_{s}}\frac{\mu_{0}\delta P_{\perp}}{B^{2}}\right].$ (16)
@@
$\delta\!P_{\perp}$ $=$ $\sum_{s}\int\mathrm{d}\mathbf{v}^{3}\,F_{0s}\left\{\frac{m_{s}v_{\perp}^{2}}{2T_{0s}}q_{s}\delta\!\phi\left[1-\frac{\omega_{\star s}^{T}}{\omega}+\frac{\omega_{ds}}{\omega}\right]\right.-\left.\frac{m_{s}^{2}v_{\perp}^{4}}{4T_{0s}}\frac{\mu_{0}\delta P_{\perp}}{B^{2}}\left[1-\frac{\omega_{\star s}^{T}}{\omega}+\frac{\omega_{ds}}{\omega}\right]\right\}$
$=$ $\underbrace{\delta\!\phi\sum_{s}\int\mathrm{d}\mathbf{v}^{3}\,\frac{q_{s}m_{s}v_{\perp}^{2}F_{0s}}{2T_{0s}}\left[1-\frac{\omega_{\star s}^{T}}{\omega}+\frac{\omega_{ds}}{\omega}\right]}_{[I]}-\underbrace{\frac{\mu_{0}\delta\!P_{\perp}}{B^{2}}\sum_{s}\int\mathrm{d}\mathbf{v}^{3}\,\frac{m_{s}^{2}v_{\perp}^{4}F_{0s}}{4T_{0s}}\left[1-\frac{\omega_{\star s}^{T}}{\omega}+\frac{\omega_{ds}}{\omega}\right]}_{[II]}$
@@
$[I]$ $=$ $\delta\!\phi\sum_{s}n_{s}q_{s}\left\{1-\frac{\omega_{\star s}}{\omega}\left(1+\eta_{s}\right)+\frac{T_{0s}}{q_{s}B\omega}\left(\omega_{\kappa}+2\omega_{\nabla B}\right)\right\}.$ (18)
The first term in (18) vanishes due to quasineutrality. On summing over species the numerator in the second term reduces to:
$\sum_{s}n_{s}q_{s}\omega_{\star s}(1+\eta_{s})=\frac{(\mathbf{k}\times\mathbf{b}\cdot\nabla{P})}{B}=\frac{B}{\mu_{0}}\left(\omega_{\kappa}-\omega_{\nabla B}\right),$ (19)
@@
$[I]$ $=\frac{B\delta\!\phi}{\mu_{0}\omega}\left(\left(\omega_{\nabla B}-\omega_{\kappa}\right)+\frac{\beta}{2}\left(\omega_{\kappa}+2\omega_{\nabla B}\right)\right)$ (20)
@@
$[II]=-\delta P_{\perp}\sum_{s}\beta_{s}\left[1-\frac{\omega_{\star s}}{\omega}(1+2\eta_{s})+\frac{T_{0s}}{q_{s}B\omega}(\omega_{\kappa}+3\omega_{\nabla B})\right]$ (21)
@@
$\delta P_{\perp}=\frac{B\delta\!\phi}{\mu_{0}\omega}\left(\omega_{\nabla B}-\omega_{\kappa}\right),$ (22)
@@
Appendix B Cancellation of Magnetic Drift and $\mathbf{\delta\!B_{\parallel}}$ Terms in Appropriate Limits
Using the relationship between $\nabla B$ and curvature drifts in 33, the drift term on the left-hand side of the GK equation in 1 can be written in a form convenient for the MHD-1 approximation:
$T_{ds}$ $=$ $-\frac{1}{\Omega_{s}}\left[\omega_{\kappa}\left(v_{\parallel}^{2}+\frac{v_{\perp}^{2}}{2}\right)-\frac{\mu_{0}(\mathbf{k}_{\perp}\cdot\mathbf{b}\times\nabla p)}{B^{2}}\frac{v_{\perp}^{2}}{2}\right]g_{s}$ (24)